# Transversal Structures in Graph Systems: A Survey

### v1

@@ 3.1 Hamilton path and Hamilton cycle
A Hamilton path (cycle) in a graph is a path (cycle) that visits every vertex exactly once. Hamiltonicity is one of the most well-studied properties in graph theory. Dirac [39] showed that every $n$-vertex graph with minimum degree at least $n/2$ contains a Hamilton cycle. Later on, Ore [117] relaxed this condition and proved that every $n$-vertex graph $G$ with $\sigma_{2}(G)\geq n$ contains a Hamilton cycle, where $\sigma_{2}(G)=\min\{d(x)+d(y):xy\notin E(G)\}$. Moon and Moser [116] established the bipartite analogues of Dirac’s theorem.
@@
The distance of two distinct vertices in a graph is the length of a shortest path connecting them. The $k$-th power $G^{k}$ of a graph $G$ is the graph obtained from $G$ by joining every pair of vertices with distance at most $k$ in $G$. Pósa (for $r=3$, [43]) and Seymour (for $r>3$, [125]) conjectured that for all $r\geq 3$ and all $n\geq r$, every $n$-vertex graph $G$ with $\delta(G)\geq{(r-1)}n/r$ must contain the $(r-1)$-th power of a Hamilton cycle. This conjecture has been confirmed by Komlós, Sárközy, and Szemerédi [90, 92]. Recently, Gupta, Hamann, Müyesser, Parczyk and Sgueglia [62] determined the asymptotically tight bound on $\delta(\mathcal{G})$ for the existence of transversal powers of Hamilton cycles, and they proposed a problem for the exact version of transversal Pósa–Seymour conjecture.
@@
Let $\mathcal{G}=\{G_{1},\ldots,G_{rn}\}$ be a graph system of order $n$. If $\delta(\mathcal{G})\geq(1-\frac{1}{r})n$, is already sufficient for a transversal copy of the $r$-th power of a Hamilton cycle?
@@
Given a graph system $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ of order $n$, what conditions imposed on $\mathcal{G}$ will guarantee that for every coloring $\chi:E(H)\rightarrow\{1,\ldots,m\}$, there is a copy of $H$ such that the image of $e$ is an edge of $G_{\chi(e)}$ for all $e\in E(H)$?
@@
Determining the exact minimum degree condition to guarantee all Hamilton cycle patterns remains an intriguing open problem. Bowtell, Morris, Pehova and Staden [19] provided a construction demonstrating that $n/2+1$ is a necessary lower bound and is conjectured as an exact threshold. In light of this, it seems more pertinent to either obtain results that improve this $o(n)$ term to an explicit and relatively small constant, or find counterexamples to this being possible.
@@
It follows from Theorem 3.5 that there exist at least $n!\sim\sqrt{2\pi n}(n/e)^{n}$ different transversal Hamilton cycles in a graph system consisting of $n$ asymptotically Dirac graphs. Bradshaw, Halasz, and Stacho [21] showed that given a system of $n$ Dirac graphs with order $n$, there are at least $(cn/e)^{cn}$ different transversal Hamilton cycles for some constant $c\geq 1/68$. Anastos and Chakraborti [11] improved the result of [21] by showing that there exists a constant $C>0$ such that there are $(Cn)^{2n}$ different transversal Hamilton cycles in a graph system consisting of $n$ Dirac graphs, which is tight up to the choice of $C$. They [11] also showed that every graph system consisting of $n$ Dirac graphs with order $n$ spans $n/2$ pairwise edge-disjoint transversal Hamilton cycles, and this is best possible.
Cheng and Staden [31] established a stability result for transversal Hamilton cycles. Given integers $a,b\geq 0$, let $\mathcal{H}_{a}^{b}$ be a graph system obtained by taking $a$ copies of $K_{\frac{n}{2}}\cup K_{\frac{n}{2}}$ and $b$ copies of $K_{{\frac{n}{2}},{\frac{n}{2}}}$ where they are defined on the same equitable partition $A\cup B$. For two vertex subsets $X,\,Y\subseteq V$, let $G[X]$ denote the subgraph of $G$ induced by $X$, let $G[X,Y]$ be the subgraph of $G$ with vertex set $X\cup Y$ and edge set $\{xy:x\in X,\,y\in Y\ \text{and}\ xy\in E(G)\}$. We say that a graph system $\mathcal{J}$ of order $n$ is half-split if there is $A\subseteq V$ with $|A|=\lfloor n/2\rfloor+1$ such that $J[A]=\emptyset$ and $J[A,V\setminus A]$ is complete for all $J\in\mathcal{J}$. Given $\kappa>0$, we say that two graph systems $\mathcal{G}$ and $\mathcal{G}^{\prime}$ on a common vertex set of size $n$ are $\kappa$-close if $\mathcal{G}^{\prime}$ can be obtained from $\mathcal{G}$ by adding or deleting at most $\kappa n^{3}$ edges.
@@
The proof of Theorem 3.6 was a combination of their newly developed transversal regularity-blow-up method [30] (see Theorem 4.4 in the next section) and the absorption method. As a continuation of the transversal version of Dirac’s theorem (Theorem 3.2) and the stability result for transversal Hamilton cycles (Theorem 3.6), Cheng, Sun, Wang and Wei [32] characterized all graph systems with order $n$ and minimum degree at least $n/2-1$, each of which does not contain transversal Hamilton cycles.
Li, Li and Li [101] studied the existence of rainbow Hamilton paths in a graph system $\mathcal{G}=\{G_{i}:i\in[n]\}$ under Ore-type conditions. However, the edges of a rainbow Hamilton path only come from $n-1$ graphs of $\mathcal{G}$. Thus, it is worth studying whether $n-1$ graphs can guarantee the existence of a rainbow Hamilton path in a graph system $\mathcal{G}=\{G_{i}:i\in[n-1]\}$. It is well-known that Dirac’s theorem implies that every $n$-vertex graph with minimum degree at least $(n-1)/2$ contains a Hamilton path. Cheng, Sun, Wang and Wei [32] proved a transversal version of this result.
@@
Similar to the case of Hamilton cycles, Cheng and Staden [31] established a stability result for transversal Hamilton paths; Cheng, Sun, Wang and Wei [32] characterized all graph systems with order $n$ and minimum degree at least $(n-3)/2$, each of which does not contain transversal Hamilton paths.
In 1971, Bondy [16] posed a meta-conjecture which states that almost all nontrivial sufficient conditions on a graph which implies that the graph is Hamiltonian also implies that the graph is pancyclic, or there might be a simple family of exceptional graphs. Some concepts similar to pancyclic have also attracted extensive attention, such as, vertex-pancyclic and panconnected. Here, we only illustrate those concepts in graph systems. Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a system of graphs on a common vertex set $V$ of order $n$. We say
@@
3.2 $F$-factor
Given two graphs $F$ and $G$, an $F$-tiling is a set of vertex-disjoint copies of $F$ in $G.$ A perfect $F$-tiling (or an $F$-factor) in $G$ is an $F$-tiling covering all vertices of $G$. Characterizing sufficient conditions for the existence of $F$-factors is one of the central areas of extremal graph theory. The most celebrated one is the Hajnal–Szemerédi theorem.
@@ 3.2 $F$-factor
Now, we survey some results on minimum degree conditions for the existence of transversal $F$-factors in graph systems. Joos and Kim [74] established an exact minimum degree condition to guarantee the existence transversal perfect matching (i.e., $K_{2}$-factor) in graph systems.
@@
Bradshaw [20] proved the analogue result in bipartite graph systems, which is equivalent to a special case of a matching theorem for $r$-partite $r$-uniform hypergraphs obtained by Aharoni, Georgakopoulos and Sprüssel [4], for the case $r=3$.
Cheng, Han, Wang and Wang [27] gave an asymptotical version of the transversal Hajnal–Szemerédi theorem. Furthermore, they proposed a conjecture for the exact version of transversal Hajnal–Szemerédi theorem.
@@
Let $\mathcal{G}=\{G_{1},\ldots,G_{\frac{n}{k}\binom{k}{2}}\}$ be a graph system of order $n$. If $\delta(\mathcal{G})\geq(1-\frac{1}{k})n$, then $\mathcal{G}$ contains transversal $K_{k}$-factor.
Montgomery, Müyesser and Pehova [115] showed that, in most cases, the same minimum degree bound is asymptotically sufficient for the existence of transversal $F$-factors, which generalized the Kühn–Osthus theorem [95] on factors.
@@
Let $\varepsilon>0$ and let $F$ be a graph on $r$ vertices with $t$ edges. If $\delta_{F{\text{-}}factor}\geq\frac{1}{2}$ or F has a bridge, then let $\delta_{F}^{T}=\delta_{F{\text{-}}factor},and$ otherwise let $\delta_{F}^{T}=\frac{1}{2}.$ Then, there is some $n_{0}$ such that the following holds for all $n\geq n_{0}.$
@@
Fischer [52] conjectured that the $k$-partite graph $G$ contains a $K_{k}$-factor if $\delta^{\prime}(G)\geq(1-1/k)n+1$ (actually his original conjecture did not include the term $+1$, but this stronger conjecture is known to be false). An approximate version of this conjecture was demonstrated independently by Keevash and Mycroft [81], and by Lo and Markström [104]. Later, Keevash and Mycroft [82] showed that the conjecture holds for any sufficiently large graph (in fact, they obtained a more general result for $r$-partite graphs with $r\geq k$). Cheng, Han, Wang and Wang [27] extended the approximate result of Fischer’s conjecture to the transversal version.
@@ 3.3 Graphs with bounded bandwidth
Montgomery, Müyesser and Pehova [115] determined asymptotically tight bound on $\delta(\mathcal{G})$ for the existence of a transversal spanning tree with maximum degree at most $cn/\log n$, generalizing the tree-embedding theorem of Komlós, Sárközy and Szemerédi [93].
It is noteworthy that many graphs, including $F$-factors, powers of Hamilton cycles, and trees, exhibit a somewhat bounded maximum degree and low connectivity. This property of low connectivity can be captured by the following notion of bandwidth. A graph $H$ has a bandwidth at most $b$ if there exists an ordering $x_{1},\cdots,x_{n}$ of $V(H)$ such that all edges $x_{i}x_{j}\in E(H)$ satisfies $|i-j|\leq b.$ Böttcher, Schacht and Taraz [18] proved the celebrated bandwidth theorem, which determines the asymptotically optimal minimum degree condition for a graph to contain a spanning subgraph $H$ with bounded maximum degree and low bandwidth.
Since the bandwidth theorem provides a common generalization of many classical results, it is natural to characterize such a generalization for the above transversal version results. Chakraborti, Im, Kim and Liu [23] proved that such a generalization holds by showing the following bandwidth theorem for graph transversals.
@@
This covers several of the above results, and is asymptotically best possible for many graphs $H$. Note that by taking each $G_{i}$ in Theorem 3.14 to be a same graph, the bandwidth theorem of Böttcher, Schacht and Taraz [18] can be deduced immediately. Therefore, Theorem 3.14 generalizes and strengthens the original bandwidth theorem.
@@ 4 Transversal via regularity
    for all triples ($(\{h,i\},j)\in\binom{[L]}{2}\times[M]$, we have that either $G_{c}^{\prime}[V_{h},V_{i}]$ = $\emptyset$ for all $c\in\mathscr{C}_{j}$, or $\mathcal{G}_{hi,j}^{\prime}:=\{G_{c}^{\prime}[V_{h},V_{i}]:c\in\mathscr{C}_{j}\}$ is $(\varepsilon,d)$-regular.
@@
Given a graph system $\mathcal{G}=\{G_{c}:c\in\mathscr{C}\}$ on $V$ and parameters $\varepsilon>0,\,d\in[0,1)$ and $L_{0}\geq 1$, the reduced graph system $\mathcal{R}=\mathcal{R}(\varepsilon,d,L_{0})$ is defined as follows. Apply Theorem 4.2 to $\mathcal{G}$ with parameters $\varepsilon,d,L_{0}$ to obtain a graph system $\mathcal{G}^{\prime}=\{G^{\prime}_{c}:c\in\mathscr{C}\}$, a partition $V_{0},\ldots,V_{L}$ of $V$ and $\mathscr{C}_{0},\ldots,\mathscr{C}_{M}$ of $\mathscr{C}$, where $V_{0},\mathscr{C}_{0}$ are the exceptional sets, $V_{1},\ldots,V_{L}$ are the vertex clusters and $\mathscr{C}_{1},\ldots,\mathscr{C}_{M}$ are the color clusters. Then =$\mathcal{R}=\{R_{1},\ldots,R_{M}\}$ is a graph system consisting of $M$ graphs each on the same vertex set $[L]$, where $hi\in R_{j}$ if and only if $\mathcal{G}_{hi,j}^{\prime}$ is $(\varepsilon,d)$-regular for each ($(\{h,i\},j)\in\binom{[L]}{2}\times[M]$.
An $n$-vertex graph $H$ is $\mu$-separable if there is $X\subseteq V(H)$ of size at most $\mu n$ such that $H-X$ consists of components of size at most $\mu n.$ With suitable small $\mu$, separable graphs includes many natural graphs, for example, $F$-factors for a fixed $F$, $2$-regular graphs, trees, powers of Hamilton cycles and graphs of small bandwidth.
@@
By utilizing the transversal blow-up lemma, Cheng and Staden [30] proved a stability result for transversal Hamilton cycles, and characterized a large class of spanning $3$-uniform linear hypergraphs $H$ such that any sufficiently large uniformly dense $n$-vertex $3$-uniform hypergraph with minimum vertex degree $\Omega(n^{2})$ contains $H$ as a subhypergraph.

### v2

@@ 3.1 Hamilton path and Hamilton cycle
A Hamilton path (cycle) in a graph is a path (cycle) that visits every vertex exactly once. Hamiltonicity is one of the most well-studied properties in graph theory. Dirac [41] showed that every graph on $n\geq 3$ vertices with minimum degree at least $n/2$ contains a Hamilton cycle. An $n$-vertex graph is a Dirac graph if its minimum degree is at least $n/2$. Later on, Ore [124] relaxed this condition and proved that every $n$-vertex graph $G$ with $\sigma_{2}(G)\geq n$ contains a Hamilton cycle, where $n\geq 3$ and $\sigma_{2}(G)=\min\{d(x)+d(y):xy\notin E(G)\}$. Moon and Moser [123] established the bipartite analogue of Dirac’s theorem.
@@
The distance of two distinct vertices in a graph is the length of a shortest path connecting them. The $k$-th power $G^{k}$ of a graph $G$ is the graph obtained from $G$ by joining every pair of vertices with distance at most $k$ in $G$. Pósa (for $r=3$, [45]) and Seymour (for $r>3$, [132]) conjectured that for all $r\geq 3$ and all $n\geq r$, every $n$-vertex graph $G$ with $\delta(G)\geq{(r-1)}n/r$ must contain the (r1)(r-1)-th power of a Hamilton cycle. This conjecture has been confirmed by Komlós, Sárközy, and Szemerédi [96, 98] for sufficiently large $n$. Recently, Gupta, Hamann, Müyesser, Parczyk and Sgueglia [65] determined the asymptotically tight bound on $\delta(\mathcal{G})$ for the existence of transversal powers of Hamilton cycles, and they proposed a problem for the exact version of transversal Pósa–Seymour conjecture.
@@
Let $\mathcal{G}=\{G_{1},\ldots,G_{(r-1)n}\}$ be a graph system of order $n$. If $\delta(\mathcal{G})\geq(1-\frac{1}{r})n$, is this already sufficient for a transversal copy of the (r1)(r-1)-th power of a Hamilton cycle?
@@
Given a graph system $\mathcal{G}=\{G_{1},\ldots,G_{m}\}$ and a graph $H$, what conditions imposed on $\mathcal{G}$ will guarantee that for every coloring $\chi:E(H)\rightarrow\{1,\ldots,m\}$, there is a copy of $H$ such that the image of $e$ is an edge of $G_{\chi(e)}$ for all $e\in E(H)$?
@@
Determining the exact minimum degree condition to guarantee all Hamilton cycle patterns remains an intriguing open problem. Bowtell, Morris, Pehova and Staden [19] provided a construction demonstrating that $n/2+1$ is a necessary lower bound and conjectured this is an exact threshold. In light of this, it seems more pertinent to either obtain results that improve this $o(n)$ term to an explicit and relatively small constant, or find counterexamples to this being possible.
@@
Bradshaw, Halasz and Stacho [21] showed that given a system of $n$ Dirac graphs with order $n$, there are at least $(cn/e)^{cn}$ different transversal Hamilton cycles for some constant $c\geq 1/68$. Theorem 3.5 recovers a stronger lower bound, that there exist at least $n!\sim\sqrt{2\pi n}(n/e)^{n}$ different transversal Hamilton cycles in a graph system $\mathcal{G}$ consisting of $n$ asymptotically Dirac graphs (i.e., $\delta(\mathcal{G})\geq(1/2+\alpha)n$ for $\alpha>0$). Anastos and Chakraborti [11] improved this by showing that there exists a constant $C>0$ such that there are $(Cn)^{2n}$ different transversal Hamilton cycles in a graph system consisting of $n$ Dirac graphs, which is tight up to the choice of $C$. They [11] also showed that every graph system consisting of $n$ Dirac graphs with order $n$ spans $n/2$ pairwise edge-disjoint transversal Hamilton cycles, and this is best possible.
Cheng and Staden [31] established a stability result for transversal Hamilton cycles. Given integers $a,b\geq 0$, let $\mathcal{H}_{a}^{b}$ be a graph system obtained by taking $a$ copies of $K_{\frac{n}{2}}\cup K_{\frac{n}{2}}$ and $b$ copies of $K_{{\frac{n}{2}},{\frac{n}{2}}}$ where they are defined on the same equitable partition $A\cup B$. For two disjoint vertex subsets $X,\,Y\subseteq V$, let $G[X]$ denote the subgraph of $G$ induced by $X$, let $G[X,Y]$ be the subgraph of $G$ with vertex set $X\cup Y$ and edge set $\{xy:x\in X,\,y\in Y\ \text{and}\ xy\in E(G)\}$. We say that a graph system $\mathcal{J}$ on the vertex set $V$ of order $n$ is half-split if there is $A\subseteq V$ with $|A|=\lfloor n/2\rfloor+1$ such that $J[A]=\emptyset$ and $J[A,V\setminus A]$ is complete for all $J\in\mathcal{J}$. Given $\kappa>0$, we say that two graph systems $\mathcal{G}$ and $\mathcal{G}^{\prime}$ on a common vertex set of order $n$ are $\kappa$-close if $\mathcal{G}^{\prime}$ can be obtained from $\mathcal{G}$ by adding or deleting at most $\kappa n^{3}$ edges.
@@
The proof of Theorem 3.6 was a combination of their newly developed transversal regularity-blow-up method [32] (see Theorem 4.4 in the next section) and the absorption method. As a continuation of the transversal version of Dirac’s theorem (Theorem 3.2) and the stability result for transversal Hamilton cycles (Theorem 3.6), Cheng, Sun, Wang and Wei [33] characterized all graph systems with order $n$ and minimum degree at least n/21n/2-1, each of which does not contain transversal Hamilton cycles.
Li, Li and Li [108] studied the existence of rainbow Hamilton paths in a graph system $\mathcal{G}=\{G_{i}:i\in[n]\}$ under Ore-type conditions. However, the edges of a rainbow Hamilton path only come from n1n-1 graphs of $\mathcal{G}$. Thus, it is worth studying whether n1n-1 graphs can guarantee the existence of a rainbow Hamilton path in a graph system $\mathcal{G}=\{G_{i}:i\in[n-1]\}$. It is well-known that Dirac’s theorem implies that every $n$-vertex graph with minimum degree at least (n1)/2(n-1)/2 contains a Hamilton path. Cheng, Sun, Wang and Wei [33] proved a transversal version of this result.
@@
Similar to the case of Hamilton cycles, Cheng and Staden [31] established a stability result for transversal Hamilton paths; Cheng, Sun, Wang and Wei [33] characterized all graph systems with order $n$ and minimum degree at least (n3)/2(n-3)/2, each of which does not contain transversal Hamilton paths.
In 1971, Bondy [16] posed a meta-conjecture which states that almost all nontrivial sufficient conditions on a graph which ensure that the graph is Hamiltonian also imply that the graph is pancyclic, or there might be a simple family of exceptional graphs. Some concepts similar to pancyclic have also attracted extensive attention, such as, vertex-pancyclic and panconnected. Here, we only illustrate those concepts in graph systems. Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a system of graphs on a common vertex set $V$ of order $n$. We say
@@
3.2 $F$-factors
Given two graphs $F$ and $G$, an $F$-tiling is a set of vertex-disjoint copies of $F$ in $G.$ A perfect $F$-tiling (or an $F$-factor) in $G$ is an $F$-tiling covering all vertices of $G$. Characterizing sufficient conditions for the existence of an $F$-factor is one of the central areas of extremal graph theory. The most celebrated one is the Hajnal–Szemerédi theorem.
@@ 3.2 $F$-factors
Now, we survey some results on minimum degree conditions for the existence of a transversal $F$-factor in graph systems. Joos and Kim [80] established an exact minimum degree condition to guarantee the existence of a transversal perfect matching (i.e., $K_{2}$-factor) in graph systems.
@@
Bradshaw [20] proved the analogous result in bipartite graph systems, which is equivalent to a special case of a matching theorem for $r$-partite $r$-uniform hypergraphs obtained by Aharoni, Georgakopoulos and Sprüssel [4], for the case $r=3$.
Cheng, Han, Wang and Wang [28] gave an asymptotic version of the transversal Hajnal–Szemerédi theorem. Furthermore, they proposed a conjecture for the exact version of transversal Hajnal–Szemerédi theorem.
@@
Let $\mathcal{G}=\{G_{1},\ldots,G_{\frac{n}{k}\binom{k}{2}}\}$ be a graph system of order $n$. If $\delta(\mathcal{G})\geq(1-\frac{1}{k})n$, then $\mathcal{G}$ contains a transversal $K_{k}$-factor.
Montgomery, Müyesser and Pehova [122] showed that, in most cases, the same minimum degree bound is asymptotically sufficient for the existence of transversal $F$-factors, which generalized the Kühn–Osthus theorem [101] on factors. For each graph $F$, let $\delta_{F}^{f}$ be the smallest real number $\delta\geq 0$ such that, for each $\epsilon>0$ there is some $n_{0}$ such that, for every $n\geq n_{0}$ with $|V(F)|$ dividing $n$, if an $n$-vertex graph $H$ has minimum degree at least $(\delta+\epsilon)n$, then $H$ contains an $F$-factor.
@@
Let $\varepsilon>0$ and let $F$ be a graph on $r$ vertices with $t$ edges. If $\delta_{F}^{f}\geq\frac{1}{2}$ or $F$ has a bridge, then let $\delta_{F}^{T}=\delta_{F}^{f}$, and otherwise let $\delta_{F}^{T}=\frac{1}{2}.$ Then, there is some $n_{0}$ such that the following holds for all $n\geq n_{0}.$
@@
Fischer [54] conjectured that the $k$-partite graph $G$ contains a $K_{k}$-factor if $\delta^{\prime}(G)\geq(1-1/k)n+1$ (actually his original conjecture did not include the term +1+1, but this stronger conjecture is known to be false). An approximate version of this conjecture was demonstrated independently by Keevash and Mycroft [87], and by Lo and Markström [111]. Later, Keevash and Mycroft [88] showed that the conjecture holds for any sufficiently large graph (in fact, they obtained a more general result for $r$-partite graphs with $r\geq k$). Cheng, Han, Wang and Wang [28] extended the approximate result of Fischer’s conjecture to the transversal version.
@@ 3.3 Graphs with bounded bandwidth
Montgomery, Müyesser and Pehova [122] determined an asymptotically tight bound on $\delta(\mathcal{G})$ for the existence of a transversal spanning tree with maximum degree at most $cn/\log n$, generalizing the tree-embedding theorem of Komlós, Sárközy and Szemerédi [99].
It is noteworthy that many graphs, including $F$-factors, powers of Hamilton cycles, and trees, exhibit a somewhat bounded maximum degree and low connectivity. This property of low connectivity can be captured by the following notion of bandwidth. A graph $H$ has a bandwidth at most $b$ if there exists an ordering $x_{1},\cdots,x_{n}$ of $V(H)$ such that all edges $x_{i}x_{j}\in E(H)$ satisfy $|i-j|\leq b.$ Böttcher, Schacht and Taraz [18] proved the celebrated bandwidth theorem, which determines the asymptotically optimal minimum degree condition for a graph to contain a spanning subgraph $H$ with bounded maximum degree and low bandwidth.
Since the bandwidth theorem provides a common generalization of many classical results, it is natural to give such a generalization for the above transversal version results. Chakraborti, Im, Kim and Liu [24] proved that such a generalization holds by showing the following bandwidth theorem for graph transversals.
@@
This covers several of the above results, and is asymptotically best possible for many graphs $H$. Note that by taking each $G_{i}$ in Theorem 3.14 to be the same graph, the bandwidth theorem of Böttcher, Schacht and Taraz [18] can be deduced immediately. Therefore, Theorem 3.14 generalizes and strengthens the original bandwidth theorem.
@@ 4 Transversal via regularity
    for all triples $(\{h,i\},j)\in\binom{[L]}{2}\times[M]$, we have that either $G_{c}^{\prime}[V_{h},V_{i}]$ = $\emptyset$ for all $c\in\mathscr{C}_{j}$, or $\mathcal{G}_{hi,j}^{\prime}:=\{G_{c}^{\prime}[V_{h},V_{i}]:c\in\mathscr{C}_{j}\}$ is $(\varepsilon,d)$-regular.
@@
Given a graph system $\mathcal{G}=\{G_{c}:c\in\mathscr{C}\}$ on $V$ and parameters $\varepsilon>0,\,d\in[0,1)$ and $L_{0}\geq 1$, the reduced graph system $\mathcal{R}=\mathcal{R}(\varepsilon,d,L_{0})$ is defined as follows. Apply Theorem 4.2 to $\mathcal{G}$ with parameters $\varepsilon,d,L_{0}$ to obtain a graph system $\mathcal{G}^{\prime}=\{G^{\prime}_{c}:c\in\mathscr{C}\}$, a partition $V_{0},\ldots,V_{L}$ of $V$ and $\mathscr{C}_{0},\ldots,\mathscr{C}_{M}$ of $\mathscr{C}$, where $V_{0},\mathscr{C}_{0}$ are the exceptional sets, $V_{1},\ldots,V_{L}$ are the vertex clusters and $\mathscr{C}_{1},\ldots,\mathscr{C}_{M}$ are the color clusters. Then $\mathcal{R}=\{R_{1},\ldots,R_{M}\}$ is a graph system consisting of $M$ graphs each on the same vertex set $[L]$, where $hi\in R_{j}$ if and only if $\mathcal{G}_{hi,j}^{\prime}$ is $(\varepsilon,d)$-regular for each $(\{h,i\},j)\in\binom{[L]}{2}\times[M]$.
An $n$-vertex graph $H$ is $\mu$-separable if there is $X\subseteq V(H)$ of size at most $\mu n$ such that HXH-X consists of components of size at most $\mu n.$ With suitable small $\mu$, separable graphs include many natural graphs, for example, $F$-factors for a fixed $F$, 22-regular graphs, trees, powers of Hamilton cycles and graphs of small bandwidth.
@@
By utilizing the transversal blow-up lemma, Cheng and Staden [32] proved a stability result for transversal Hamilton cycles, and characterized a large class of spanning $3$-uniform linear hypergraphs $H$ such that any sufficiently large uniformly dense $n$-vertex 33-uniform hypergraph with minimum vertex degree $\Omega(n^{2})$ contains $H$ as a subhypergraph.